# Self-Organising Memristive Networks as Physical Learning Systems

### v1

@@ 4.2 Associative learning
Associative learning, i.e. the formation of associations between stimuli and responses, can be implemented in SOMNs by taking advantage of their adaptive conductance changes (i.e. plasticity, cf. sec. 2.2). External electrical stimuli applied to SOMNs drive local conductance changes that can lead to global reconfigurations in network states as unique conductance pathways are established for different patterns in the stimuli [161, 46]. The memory encoded in their conductance states [56] enables SOMNs to recall, or self–learn, these input–output associations. In contrast to reservoir computing, learning is intrinsic to the network: repeated stimulation leads to material-level changes in conductance pathways, enabling SOMNs to self-organise and recall associations without external training mechanisms. This is more analogous to biological synaptic plasticity than statistical optimisation. Teaching SOMNs input–output relationships via feedback, without requiring explicit programming also contrasts sharply with conventional implementations of physical neural networks, which typically rely on direct internal state addressability, for example, to physically implement Hopfield associative memory networks [162].
@@
Both positive and negative feedback enables SOMNs to adapt their electrical connectivity and conductance paths iteratively, to promote the emergence of robust functional mappings. This iterative correction process can be viewed as a form of evolutionary search, in which input–output feedback functions as a fitness signal, guiding the system toward functional configurations [165, 166]. In this view, learning resembles an adaptive exploration of the network’s dynamical state space, governed by error correction, constrained by physical laws and driven by environmental cues [37]. These mechanisms underscore that learning is not defined by task complexity, but rather by the system dynamics and efficiency of feedback. SOMNs bridge conventional and biological models, offering real–time adaptability, dynamic reconfiguration, and connectivity–driven behaviour [35, 37]. A key open question is how many states can be learned and/or stored in these types of physical architectures?
@@ 5 Outlook
The brain demonstrates that intelligence can emerge from the physical properties of matter. This realisation continues to motivate the search for materials and architectures with innate brain-like physical properties that may thus be considered promising substrates for emulating core cognitive functions such as memory and learning. Self-organising memristive networks are compelling candidates in this endeavour. As highlighted in this Perspective, the emergent nonlinear dynamics produced by these non-equilibrium physical systems can be exploited in adaptive learning paradigms, without the need for explicit programming. This sets the stage for developing broader technological applications and exploring long-term implications.
@@
Central to their role in physical learning systems is the emergence of global, collective dynamics from relatively simple, local interactions. Nanoscale transport processes (i.e. electron and ionic migration, and atomic rearrangement) continuously modulate junction-level conductance. These local dynamics propagate globally under Kirchhoff constraints and the complex network topology gives rise to dynamical phase transitions, a hallmark of emergent phenomena. This interplay between local plasticity and global self-organisation underpins the functional properties that serve to encode information and enable learning.
@@
Here, we have highlighted two learning paradigms, reservoir computing and associative learning, that leverage the internal physics of self–organising memristive networks. A range of other schemes are now under development, such as those based on spiking computation[82], probabilistic computing [83, 84] and energy–based algorithms (e.g. Hopfield networks, Boltzmann machines) that use binary units which may be physically instantiated with thresholded memristive switching. Recently it was shown that SOMNs that incorporate both spiking neural characteristics and memristive synapses allow new types of learning behaviour [167]. Another research question that arises is: what determines the upper bound of the ability of these memristive networks to process information (i.e. compression, coding/decoding of data)? In some cases, it may be possible to minimise the amount of external computation required to obtain a desired output, effectively making the learning training–free and independent of learning algorithms [168].
From a theoretical standpoint, progress in modeling memristive networks, using tools such as graph-theoretic projector operators or mean-field approximations, continues to enhance our understanding of their functional capacity. These models enable us to describe and even guide the evolution of useful dynamical states, allowing for a form of guided self-organisation [169]. By tuning network density, topology, or stimulation protocols, one can steer the system toward regimes that balance flexibility, stability, and computational richness. On a pragmatic level, recent modelling work has demonstrated the way that different memristor types affect reservoir computing performance [170]
Altogether, self-organising memristive networks represent a material route toward physically embodied intelligence. They shift the paradigm from computing architectures that are engineered and explicitly programmed, to systems that are grown and trained through physical interaction and feedback. By embedding function into structure, and adaptation into dynamics, these materials may offer a path toward scalable, autonomous learning platforms that mirror both the form and function of biological cognition. Taken together, this convergence of physical dynamics and computational function positions SOMNs not merely as neuromorphic analogs, but as a distinct, material foundation for intelligent physical systems.
Ultimately, such systems suggest a future where intelligence is not imposed onto matter, but emerges from it; where computation is not merely symbolic, but physical; not strictly deterministic, but adaptive; and not rigidly encoded, but continuously evolving. In this framework, self–organising memristive networks are a resource that can be harnessed for a new generation of brain–inspired physical learning architectures.
@@
Acknowledgements. F.C.’s work was conducted under the auspices of the National Nuclear Security Administration of the United States Department of Energy at Los Alamos National Laboratory (LANL) under Contract No. DE-AC52-06NA25396, and supported financially via DOE LDRD grant 20240245ER. S.A.B acknowledges support from the MacDiarmid Institute for Advanced Materials and Nanotechnology and the Marsden Fund NZ. G.M and C.R. acknowledge funding by NEURONE, a project funded by the European Union - Next Generation EU, M4C1 CUP I53D23003600006, under program PRIN 2022 (prj code 20229JRTZA). G.M. acknowledges funding by the European Union (ERC, "MEMBRAIN", No. 101160604). Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.

### v2

@@ 4.2 Associative learning
Associative learning, i.e. the formation of associations between stimuli and responses, can be implemented in SOMNs by taking advantage of their adaptive conductance changes (i.e. plasticity, cf. sec. 2.2). External electrical stimuli applied to SOMNs drive local conductance changes that can lead to global reconfigurations in network states as unique conductance pathways are established for different patterns in the stimuli [183, 53]. The memory encoded in their conductance states [62] enables SOMNs to recall, or self–learn, these input–output associations. In contrast to reservoir computing, where learning occurs in an external layer, associative learning occurs directly inside the SOMN via adjustment of junction conductance states. As repeated stimulation leads to material-level changes in conductance pathways, SOMNs self-organise and recall associations between input patterns and global conductance states. This is more analogous to biological synaptic plasticity than statistical optimisation. Teaching SOMNs input–output relationships via feedback, without requiring explicit programming also contrasts sharply with conventional implementations of physical neural networks, which typically rely on direct internal state addressability, for example, to physically implement Hopfield associative memory networks [184].
@@
Both positive and negative feedback enable SOMNs to iteratively adapt their electrical connectivity and conductance pathways, thereby promoting the emergence of robust functional mappings. This iterative correction process can be viewed, by analogy with genetic algorithms, as a form of evolutionary search in which the agreement between the actual output and the desired target acts as a fitness signal: configurations that reduce the output error are effectively reinforced, whereas less successful ones are suppressed or redirected [187, 188]. In this sense, learning can be interpreted as an adaptive exploration of the network’s dynamical state space, driven by input–output feedback, constrained by physical laws, and shaped by environmental cues [37]. These mechanisms underscore that learning is not defined by task complexity, but rather by the system dynamics and efficiency of feedback. SOMNs bridge conventional and biological models, offering real–time adaptability, dynamic reconfiguration, and connectivity–driven behaviour [35, 37]. A key open question is how many states can be learned and/or stored in these types of physical architectures?
@@ 5 Outlook
The brain demonstrates that intelligence can emerge from the physical properties of matter. This realisation continues to motivate the search for materials and architectures with innate brain-like physical properties that may thus be considered promising substrates for emulating core cognitive functions such as memory and learning. Self-organising memristive networks are compelling candidates in this endeavour. As highlighted in this Perspective, the emergent nonlinear dynamics produced by these non-equilibrium physical systems can be exploited in adaptive learning paradigms, without the need for explicit programming.
This sets the stage for developing broader technological applications and exploring long-term implications.
@@
Central to their role in physical learning systems is the emergence of global, collective dynamics from relatively simple, local interactions. Nanoscale transport processes (i.e. electron and ionic migration, and atomic rearrangement) continuously modulate junction-level conductance. These local dynamics propagate globally under Kirchhoff constraints and the complex network topology, giving rise to dynamical phase transitions, a hallmark of emergent phenomena. This interplay between local plasticity and global self-organisation underpins the functional properties that serve to encode information and enable learning.
A key opportunity for future work is to move SOMNs from intriguing functional materials toward engineerable learning substrates. We anticipate that this will require physics-grounded modelling pipelines that can begin to connect controllable fabrication parameters and stimulation protocols to task-level performance, although much remains to be understood before such links can be made predictive in practice. Hierarchical and multi-scale approaches, ranging from junction-level models and Kirchhoff-constrained network dynamics to reduced descriptions such as mean-field and stochastic theories, offer a promising route toward this goal. In principle, such approaches could help identify favourable operating regimes—for example in terms of memory depth, nonlinearity, criticality, stability, stimulation strategies, and robustness–energy trade-offs—while also clarifying the collective non-equilibrium physics of disordered adaptive circuits. Establishing these connections is therefore both a fundamental scientific challenge and an exciting practical step toward deploying SOMNs for embedded and continuous learning. In parallel, Equilibrium Propagation, generalized Equilibrium Propagation, and coupled physical learning suggest plausible routes toward in-hardware training, potentially even beyond the quasistatic limit, although substantial theoretical and experimental work is still needed to determine under what conditions such approaches can be implemented reliably and how bias or finite-nudge effects can be controlled [189, 190, 191, 192, 193].
@@
Here, we highlight two learning paradigms, reservoir computing and associative learning, that leverage the intrinsic physical dynamics of self-organising memristive networks. Other schemes that are now under development include those based on spiking computation[92], probabilistic computing [93, 94] and energy–based algorithms (e.g. Hopfield networks, Boltzmann machines) that use binary units which may be physically instantiated with thresholded memristive switching. SOMNs that incorporate both spiking neural characteristics and memristive synapses can allow new types of learning behaviour [104]. The incorporation of molecular synapses into SOMNs is an exciting new avenue for research as the introduction of additional components into the networks (beyond those described in Section 2) provides new flexibility to tune network properties and opens up myriad new possibilities to implement new styles of computation.
Alongside these emerging computational schemes, an important theoretical frontier is to extend the present modelling and learning framework from predominantly two-terminal descriptions to genuinely multi-terminal operation. This extension is especially timely because many of the new functionalities discussed above—including richer spiking behaviour, hybrid synaptic mechanisms, and more flexible forms of computation—are most naturally expressed in architectures that can be driven and read out through several electrodes simultaneously. In realistic experimental SOMNs, multi-terminal configurations enable distributed readout, spatially structured stimulation, and more complex boundary conditions than those captured by an effective two-point conductance description. Developing Kirchhoff-constrained network dynamics, reduced models, and learning rules for such multi-point contacts is therefore a key opportunity for future work, both to broaden the range of accessible tasks and to provide practical design principles for electrode placement, stimulation protocols, and scalability in deployable devices.
A further opportunity is to understand what ultimately sets the upper bound on the information-processing capacity of these memristive networks, including their ability to compress, transform, and encode data. Addressing this question would not only clarify the computational limits of SOMNs, but could also reveal regimes in which the physical substrate itself performs a larger share of the computation. In favourable cases, this may reduce the amount of external training or post-processing required to obtain a desired output, raising the exciting possibility of partially training-free learning schemes that rely more directly on the intrinsic dynamics of the material system [194].
More broadly, SOMNs open a rich design space in which physical structure, adaptive dynamics, and task performance may be co-optimized rather than treated separately. Understanding how network density, topology, material composition, and stimulation protocols shape learning behaviour remains an important challenge, but also a major opportunity for future work. In this respect, SOMNs provide a natural meeting point between neuromorphic engineering, statistical physics, and the complex systems community, where concepts such as emergence, adaptability, and collective computation can be brought to bear on the design of physical learning substrates.
This perspective also connects naturally to the broader idea of guided self-organization [136]: rather than prescribing function entirely through explicit programming, one seeks to steer a physical system toward useful dynamical regimes through appropriate constraints, feedback, and environmental coupling. For SOMNs, this suggests an appealing long-term vision in which learning is not simply imposed on the hardware, but arises from the controlled interplay between material properties, network architecture, and stimulation. Developing such principles would deepen our understanding of self-organizing adaptive matter while also providing practical routes toward more autonomous and energy-efficient physical learning systems. Taken together, these directions suggest that self-organising memristive networks may become more than intriguing neuromorphic analogs. They offer the possibility of a distinct material foundation for intelligent physical systems, in which computation, memory, and adaptation are co-localised in the same substrate. In this future-looking view, SOMNs are not only a platform for implementing brain-inspired ideas, but also an opportunity to discover new forms of computation emerging from disordered, adaptive, and self-organizing matter.
@@
Acknowledgements. F.C.’s work was conducted under the auspices of the National Nuclear Security Administration of the United States Department of Energy at Los Alamos National Laboratory (LANL) under Contract No. DE-AC52-06NA25396, and supported financially via DOE LDRD grant 20240245ER. F.C. is now an employee of Planckian. S.A.B acknowledges support from the MacDiarmid Institute for Advanced Materials and Nanotechnology and the Marsden Fund NZ. G.M and C.R. acknowledge funding by NEURONE, a project funded by the European Union - Next Generation EU, M4C1 CUP I53D23003600006, under program PRIN 2022 (prj code 20229JRTZA). G.M. acknowledges funding by the European Union (ERC, "MEMBRAIN", No. 101160604). Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.